OpenStudy (anonymous):

4. 3s+6<-5(s+2) plz help

OpenStudy (anonymous):

3s+6<-5(s+2) first we do -5(s+2)=-5 times s and -5 times 2 = -5s -10 3s+6<-5s-10 = 3s+5s<-6-10 (we flip to make variable with variable and number with number)= 8s<-16 for get the s we should do -16 divided 8, s = -2 for verification : 3(-2)+6<-5(-2+2) = -6+6 < 10-10 = 0<0